suppose that you plan to buy an apartment 7 yeara from today. you plan to save $5000 per year beginning immediately. your interest rate is 4%. how much will you have after 7 years?
Answer #1

Annuity due = 5000
Rate = 4%
Number of years = 7
FV using annuity due formula = (1+r)* Annuity Due *(1-(1+r)-n)/r = (1+4%)*5000*(1-(1+4%)-7)/4% = 31,210.68

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well